How far apart are two electrons if they exert a repulsive force of 2 N on each other?
Answer #1
F= K *Q*q/r^2 if F = 2N K=9*10^9 Q=q=-1.602*10^-19 C
r^2= K*Q^2/F = 9*10^9*(-1.602*10^-19)^2 /2= 1.15*10^-28
r= 1.075*10^-14m